OpenStudy (anonymous):

Factor and solve a^3n -b^3n

OpenStudy (kanwal32):

is it \[a ^{3n}\]

OpenStudy (kanwal32):

a^n-b^n(a^2n+b^2n+a^nb^n)

OpenStudy (anonymous):

I'm completely lost....... I don't understand :( sorry

OpenStudy (kanwal32):

ok we know that a^3-b^3=(a-b)(a^2+b^2+ab)

OpenStudy (kanwal32):

instead of a we are using a^n

OpenStudy (anonymous):

So I use sum or difference of cubes?

OpenStudy (kanwal32):

difference of cubes

OpenStudy (kanwal32):

also if (a^n)^2=a^2n we multiply the powers

OpenStudy (anonymous):

\[a^n-b^n(a^2n+b^2n+a^nb^n)\] this general idea, just to make it clear is not like this \[a^n-b^n(a2n+b2n+anbn)\]

OpenStudy (anonymous):

I got the same answer as you @Andresfon12

OpenStudy (kanwal32):

@Maddiem is it \[a ^{3n}\]

OpenStudy (anonymous):

yes

OpenStudy (kanwal32):

so as i have done it above is the right procedure

OpenStudy (anonymous):

that just a general formula he @ kanwal32 gave you, no the answer

OpenStudy (anonymous):

OKay! thanks guys for the help :D

OpenStudy (anonymous):

np

OpenStudy (kanwal32):

welcome